OpenStudy (anonymous):

Given the complex number 3 + 3i, what is theta in the polar form?

OpenStudy (helder_edwin):

\[ \theta=\arctan(3/3) \]

OpenStudy (anonymous):

\[\huge \color{green}{\theta = \tan^{-1}\frac{b}{a}}\]

OpenStudy (anonymous):

Here, b = 3 and a = 3 So what you get??

OpenStudy (anonymous):

45

OpenStudy (anonymous):

Yes you are right..

OpenStudy (anonymous):

awesome thanks!

OpenStudy (anonymous):

45 is in degrees.. If you want your answer in radians, then multiply it by: \[\pi /180\]

OpenStudy (anonymous):

pi/4

OpenStudy (anonymous):

thanks so helpful!!!!!!

OpenStudy (anonymous):

Yes absolutely..